OpenStudy (anonymous):

a 25 foot ladder leans against a building. The angle of elevation of the ladder to the building is 70 degrees. How his is the top of the ladder on the building?

OpenStudy (akashdeepdeb):

 Use trigonometry. As 25 ft is the length of the ladder, which is the hypotenuse. And we need to find the base which is 'x' meters in length. We use the formula of: \[\cos \theta = \frac{base}{hypotenuse} = \cos 70 = \frac{x}{25}\] Now can you find, how high the top of the ladder is? :D